### v1

@@ 11 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22
$S=\left\{i=2,\ldots,p_{2}:i=\left\lceil(j-C_{c}+\left\lfloor C_{c}\right\rfloor)\frac{p_{2}}{b}\right\rceil+1\mbox{, for }j=1,2,\ldots,b\right\}.$
Proof.
The formulas for $k_{1}$ and $k_{p_{2}+1}$ follow directly by substituting $p_{1}=ap_{2}+b$ into (4). Doing the same for the remaining $k_{i}$, we obtain for $i=2,\ldots,p_{2}$,
$k_{i}$ $=\left\lfloor\frac{p_{1}}{p_{2}}(i-1)+C_{c}\right\rfloor-\left\lfloor\frac{p_{1}}{p_{2}}(i-2)+C_{c}\right\rfloor$
$=\left\lfloor a(i-1)+\frac{b}{p_{2}}(i-1)+C_{c}\right\rfloor-\left\lfloor a(i-2)+\frac{b}{p_{2}}(i-2)+C_{c}\right\rfloor$
$=a+\left\lfloor\frac{b}{p_{2}}(i-1)+C_{c}\right\rfloor-\left\lfloor\frac{b}{p_{2}}(i-2)+C_{c}\right\rfloor.$
Since $\frac{b}{p_{2}}<1$, these exponents are equal to either $a$ or $a+1$. We claim at most $b$ of the $k_{i}$ are equal to $a+1$. To see this, we substitute $i^{\prime}=p_{2}$, $i=1$ and $p_{1}=ap_{2}+b$ into (2), to obtain
Finally, to determine which of the $k_{i}$ is equal to $a+1$, consider the quantity $K_{i}=\lfloor(i-1)\frac{b}{p_{2}}+C_{c}\rfloor$. If $i=1$, $K_{i}=\lfloor C_{c}\rfloor$. As $i$ increases, the values of $K_{i}$ are constant until some smallest $i$ (call it $i_{1}$) satisfies $K_{i_{1}}=\lfloor C_{c}\rfloor+1$; thus $i_{1}$ is the smallest index such that $k_{i}=a+1.$ Similarly, if $i_{2}$ is the smallest $i$ such that $K_{i_{2}}=\lfloor C_{c}\rfloor+2$ then $i_{2}$ is the second smallest index such that $k_{i}=a+1.$ More generally, the indices for which $k_{i}=a+1$ occur at the smallest indices $i_{j}$ for which $K_{i_{j}}=\lfloor C_{c}\rfloor+j.$
Let $\ell_{j}=\lceil(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}\rceil+1$; we claim $i_{j}=\ell_{j}$ for $j\geq 1$. To see this, note that
$(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}\leq\ell_{j}-1<(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}+1,$
which implies
$j+\lfloor C_{c}\rfloor\leq(\ell_{j}-1)\frac{b}{p_{2}}+C_{c}<j+\lfloor C_{c}\rfloor+\frac{b}{p_{2}},$
and hence $\lfloor(\ell_{j}-1)\frac{b}{p_{2}}+C_{c}\rfloor=j+\lfloor C_{c}\rfloor.$
On the other hand
$(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}-1\leq\ell_{j}-2<(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b},$
which implies
$j+\lfloor C_{c}\rfloor-\frac{b}{p_{2}}\leq(\ell_{j}-2)\frac{b}{p_{2}}+C_{c}<j+\lfloor C_{c}\rfloor,$
and so $\lfloor(\ell_{j}-2)\frac{b}{p_{2}}+C_{c}\rfloor<j+\lfloor C_{c}\rfloor.$ Thus, $\ell_{j}$ is the smallest integer such that $K_{\ell_{j}}=\lfloor C_{c}\rfloor+j$ or $i_{j}=\ell_{j}$, as required. ∎
@@
Suppose $p_{1}=p_{2}+1$ for some $p_{2}>1$. Then the apportionment sequence is
[…]
If $p_{2}=1$ and $p_{1}=2$, then the apportionment sequence for all stationary methods is $1^{2},2$.
@@
Suppose $p_{1}=p_{2}+1$ for some $p_{2}>1$. Then $C_{c}=(1-1+\frac{1}{p_{2}})c=\frac{c}{p_{2}}$ and $\lfloor C_{c}\rfloor=0$ for all $c\in[0,1]$. Thus, $k_{1}=1$. Moreover, if $c<1$, then (from the definition of the set $S$) we have $\lceil(1-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{1}\rceil+1=\lceil(1-\frac{c}{p_{2}})p_{2}\rceil+1=p_{2}+1$, and so $S$ is empty. Thus, $k_{2}=\cdots=k_{p_{2}}=1$, leading to the apportionment sequence 1, 2, 1, 2,, $1,\,2,\ldots,\,1$. If $c=1$ then $\lceil(1-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{1}\rceil+1=\lceil(1-\frac{c}{p_{2}})p_{2}\rceil+1=p_{2}$, which implies $S=\{p_{2}\}$. Thus $k_{2}=\cdots=k_{p_{2}-1}=1$ and $k_{p_{2}}=2$, leading to the apportionment sequence 1, 2, 1, 2,, $1,\,2,\ldots,\,1^{2},\,2$. The last statement of the proposition follows similarly. ∎
When $p_{1}>p_{2}+1$, we obtain the following.
@@
Suppose $p_{1}$ and $p_{2}$ are relatively prime and $p_{1}>p_{2}+1$. Then the apportionment sequences corresponding to stationary divisor methods are constant for cut points in the intervals $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ where $\ell=1,\ldots,p_{1}-p_{2}-1$. Moreover, each of these $p_{1}-p_{2}-1$ sequences is distinct.
Proof.
Suppose $p_{1}=ap_{2}+b$ for some integers $a\geq 0$ and $0<b<p_{2}$ and that $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ for some $\ell=1,\ldots,p_{1}-p_{2}-1$. Let $\ell=\nu p_{2}+\ell^{\prime}$ for some integers $\nu\geq 0$ and $0\leq\ell^{\prime}<p_{2}.$ Then $C_{c}=\frac{p_{1}-p_{2}}{p_{2}}c$ satisfies $\frac{\ell}{p_{2}}\leq C_{c}<\frac{\ell+1}{p_{2}}$, which implies $\lfloor C_{c}\rfloor=\nu$ and $k_{1}=\nu+1$. Thus, $k_{1}$ is constant on this interval.
To show that the other $k_{i}$ are constant on $\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ it suffices to show that the indices $i_{j}=\lceil(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}\rceil+1$ are also constant on this interval. But
$\frac{-\ell^{\prime}-1}{p_{2}}<-C_{c}+\lfloor C_{c}\rfloor\leq\frac{-\ell^{\prime}}{p_{2}},$
which implies
$\frac{jp_{2}-\ell^{\prime}-1}{b}<(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}\leq\frac{jp_{2}-\ell^{\prime}}{b}.$
So $i_{j}=\lceil(j+\lfloor C_{c}\rfloor-C_{c})\frac{p_{2}}{b}\rceil+1=\lceil\frac{jp_{2}-\ell^{\prime}}{b}\rceil+1$ is constant on the interval as claimed.
Suppose that $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ and $c^{\prime}\in[\frac{\ell+1}{p_{1}-p_{2}},\frac{\ell+2}{p_{1}-p_{2}})$ and that the two apportionment sequences are the same. Then $C_{c^{\prime}}=\frac{p_{1}-p_{2}}{p_{2}}c^{\prime}$ satisfies $\frac{\ell+1}{p_{2}}\leq C_{c^{\prime}}<\frac{\ell+2}{p_{2}}$, and recalling that $\ell=\nu p_{2}+\ell^{\prime},$ we obtain $\nu+\frac{\ell^{\prime}+1}{p_{2}}\leq C_{c^{\prime}}<\nu+\frac{\ell^{\prime}+2}{p_{2}}.$ Since the $k_{1}$ values for each sequence are equal, $\lfloor C_{c^{\prime}}\rfloor=\lfloor C_{c}\rfloor=\nu$. Hence $\ell^{\prime}+1<p_{2}$.
$\left\lceil\frac{jp_{2}-\ell^{\prime}}{b}\right\rceil=\left\lceil(j-C_{c}+\lfloor C_{c}\rfloor)\frac{p_{2}}{b}\right\rceil=\left\lceil(j-C_{c^{\prime}}+\lfloor C_{c^{\prime}}\rfloor)\frac{p_{2}}{b}\right\rceil=\left\lceil\frac{jp_{2}-(\ell^{\prime}+1)}{b}\right\rceil.$
Since the apportionment is constant on each interval of the form $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$, we can substitute $c=\frac{\ell}{p_{1}-p_{2}}$ into the expressions in Theorem 9 to get the following.
@@
Suppose $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}>p_{2}+1$ and $p_{1}=ap_{2}+b$ for some integers $a\geq 0$ and $0<b<p_{2}$. If $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ with $\ell=\nu p_{2}+\ell^{\prime}$ then for $i=2,\ldots,p_{2}$,
@@
$k_{p_{2}+1}$ $=p_{1}-(k_{1}+\ldots+k_{p_{2}}).$
Corollary 12 implies that the values of $k_{2},\ldots,k_{p_{2}}$ depend only on $\ell^{\prime}$ (and $b$), and thus are do not change for different values of $k_{1}=\nu+1$. We illustrate this in the following example.
@@
Suppose $p_{2}=7$ and $p_{1}=7a+2$ for some integer $a$. Then for each value of $k_{1}$, there are 7 different apportionment sequences corresponding to $c\in[\frac{\ell}{p_{1}-7},\frac{\ell+1}{p_{1}-7})$ where $\ell=\nu 7+\ell^{\prime}=(k_{1}-1)7+\ell^{\prime}$, as shown in Table 2. For each $\ell^{\prime}=0,\ldots 6$, the position of the exponents for which $k_{i}=a+1$ can be determined by evaluating $i_{j}=\lceil\frac{7j-\ell^{\prime}}{2}\rceil+1.$ Thus, when $\ell^{\prime}=0$ (the first row in the table), this occurs when $i=\lceil\frac{7(1)-0}{2}\rceil+1=5$. When $\ell^{\prime}=1$, this occurs when $i=\lceil\frac{7(1)-1}{2}\rceil+1=4$, and so on.
@@
Table 1 also include a row corresponding to $c=1$ (D’Hondt’s method). It is the only sequence that does not end in a “1.” In fact, this is true more generally.
@@
Suppose $p_{1}$ and $p_{2}$ are relatively prime and the apportionment sequence corresponding to the stationary divisor method with cut point $c$ is
$1^{k_{1}}\;2\;1^{k_{2}}\;2\;\ldots 1^{k_{j}}\;2\;\ldots\;1^{k_{p_{2}}}\;2\;1^{p_{1}-(k_{1}+\cdots+k_{p_{2}})}.$
@@
Substituting $i=p_{2}$ into (5), we get
@@
If $c=1$ then $\frac{p_{2}}{p_{2}-1+c}=1,$ which implies
@@
Hence $\sum_{j\leq p_{2}}k_{j}\leq p_{1}<\sum_{j\leq p_{2}}k_{j}+1,$ which implies $p_{1}=\sum_{j\leq p_{2}}k_{j}.$
Alternatively, if $p_{1}=\sum_{j\leq p_{2}}k_{j}$, then the left-hand inequality in 7) implies
@@
Note that if $c=1$ and $p_{1}=k_{1}+\cdots+k_{p_{2}}$ then the left-hand inequality in 7 is an equality: $\frac{p_{1}}{p_{1}-1+1}=\frac{p_{2}}{p_{2}-1+1}=1.$ Thus, the last two seats for each party are awarded, first to $p_{1}$ and then to $p_{2}$ using the tie-breaking rule.
@@
Proposition 14.
If $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}=ap_{2}+b$, then the apportionment sequence corresponding to D’Hondt’s method is
@@
where $k_{i}=a+1$ if $i\in S$ and $k_{i}=a$ otherwise, and $S=\{\lceil\frac{p_{2}}{b}\rceil,\lceil\frac{2p_{2}}{b}\rceil,\ldots,\lceil p_{2}\rceil\}.$
@@
If $p_{1}$ and $p_{2}$ are relatively prime then there are exactly $p_{1}-p_{2}+1$ different apportionment sequences corresponding to stationary divisor methods.
@@
Suppose $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}>p_{2}+1$ and the apportionment sequence has the form
@@
where the $k_{i}$ satisfy (1) and (2). Then the apportionment sequence corresponds to a stationary divisor method with cut point $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}+p_{2}})$ where $\ell=\nu p_{2}+\ell^{\prime}$, $\nu=k_{1}-1$ and $\ell^{\prime}=\max_{i_{j}\in S}\{p_{2}j-(i_{j}-1)b\}$ where $S=\{i_{j},|k_{i_{j}}=\lfloor\frac{p_{1}}{p_{2}}\rfloor+1,2\leq i_{j}\leq p_{2}\}$.
We conclude this by remarking that all the results of this section apply to more general values of $\mathbf{p}=(p_{1},p_{2})$ with $p_{1}$ and $p_{2}$ replaced by $p_{1}/\gcd(p_{1},p_{2})$ and $p_{2}/\gcd(p_{1},p_{2})$ in all expressions for $k_{i}$, etc.

### v2

@@ 11 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22
$S=\left\{i=2,\ldots,p_{2}:i=\left\lceil(j-\delta_{c}+\left\lfloor\delta_{c}\right\rfloor)\frac{p_{2}}{b}\right\rceil+1\mbox{, for }j=1,2,\ldots,b\right\}.$
We illustrate Theorem 3 for $(p_{1},p_{2})=(16,7)$ and two different values of $c$.
Example 6.
Let $(p_{1},p_{2})=(16,7)$. Since $16=2\cdot 7+2$, then $\delta_{c}=\left(2-1+\frac{2}{7}\right)c=\frac{9}{7}c$ and
$S=\left\{i=2,\ldots,7:i=\left\lceil\left(j-\textstyle\frac{9}{7}c+\left\lfloor\textstyle\frac{9}{7}c\right\rfloor\right)\textstyle\frac{7}{2}\right\rceil+1\mbox{, for }j=1,2\right\}.$
If $c=0$, then $\delta_{0}=\lfloor\delta_{0}\rfloor=0$. Moreover, since
$\left\lceil(1-0+\lfloor 0\rfloor)\textstyle\frac{7}{2}\right\rceil+1=5\quad\mbox{and}\quad\left\lceil(2-0+\lfloor 0\rfloor)\textstyle\frac{7}{2}\right\rceil+1=8>7,$
then $S=\{5\}$. Applying the formulas in Theorem 3, we obtain
$k_{1}=\lfloor 0\rfloor+1=1,\quad k_{5}=3,\quad k_{2},k_{3},k_{4},k_{6},k_{7}=2,\quad\mbox{ and }\quad k_{8}=3-\left\lfloor\textstyle\frac{12}{7}+0\right\rfloor=2.$
This is the sequence that appears in the first line of Table 1.
If $c=\frac{6}{9}$ then $\delta_{\frac{6}{9}}=\frac{6}{7}$ and $\lfloor\delta_{\frac{6}{9}}\rfloor=0$. Additionally,
$\left\lceil\left(1-\textstyle\frac{6}{7}+0\rfloor\right)\textstyle\frac{7}{2}\right\rceil+1=2\quad\mbox{and}\quad\left\lceil\left(2-\textstyle\frac{6}{7}+\lfloor 0\rfloor\right)\textstyle\frac{7}{2}\right\rceil+1=5,$
so $S=\{2,5\}$ and
$k_{1}=\lfloor 0\rfloor+1=1,\quad k_{2},k_{5}=3,\quad k_{3},k_{4},k_{6},k_{7}=2,\quad\mbox{ and }\quad k_{8}=3-\left\lfloor\textstyle\frac{12}{7}+\textstyle\frac{6}{7}\right\rfloor=1.$
This sequence appears in line 7 of Table 1.
@@
If $p_{1}=p_{2}+1$ and $c\in[0,1)$, then the apportionment sequence under the stationary divisor method with cutpoint $c$ has repeated pattern
1, $\underbrace{1,\,2,\,1,\,2,\,\ldots,}_{p_{2}\text{ times}}$1.
@@
Suppose $p_{1}=p_{2}+1$ for some $p_{2}\geq 1$ and $c\in[0,1)$. From Theorem 3, $\delta_{c}=(1-1+\frac{1}{p_{2}})c=\frac{c}{p_{2}}$ and $\lfloor\delta_{c}\rfloor=0$, implying $k_{1}=1$. Moreover, substituting $j=1$ into the definition of the set $S$, we have $\lceil(1-\delta_{c}+\lfloor\delta_{c}\rfloor)\frac{p_{2}}{1}\rceil+1=\lceil(1-\frac{c}{p_{2}})p_{2}\rceil+1=p_{2}+1$, and so $S=\emptyset.$ Thus, $k_{2}=\cdots=k_{p_{2}}=1$ and $k_{p_{2}+1}=1-\lfloor\frac{1}{p_{2}}(p_{2}-1)+\frac{c}{p_{2}}\rfloor=1$, leading to the repeated pattern 1, 2, 1, 2,, $1,\,2,\ldots,\,1$. ∎
When $p_{1}>p_{2}+1$, Theorem 3 can be used to determine subintervals for the values of $c\in[0,1]$ for which the apportionment sequence is constant. As suggested by the patterns in Table 1, these occur on intervals of length $1/(p_{1}-p_{2})$. This is described in the next theorem, the proof of which appears in the appendix.
@@
Suppose $p_{1}$ and $p_{2}$ are relatively prime and $p_{1}>p_{2}+1$. Then the apportionment sequences corresponding to stationary divisor methods are constant for cutpoints in the intervals $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ where $\ell=0,\ldots,p_{1}-p_{2}-1$. Moreover, each of these $p_{1}-p_{2}$ sequences is distinct.
Since the apportionment is constant on each interval of the form $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$, we can substitute $c=\frac{\ell}{p_{1}-p_{2}}$ into the expressions in Theorem 3 to get the following simplified expressions for the $k_{i}$ values.
@@
Suppose $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}>p_{2}+1$ and $p_{1}=ap_{2}+b$ for some integers $a\geq 1$ and $0<b<p_{2}$. If $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}})$ with $\ell=\nu p_{2}+\ell^{\prime}$, then, for $i=2,\ldots,p_{2}$,
@@
$k_{p_{2}+1}$ $=p_{1}-(k_{1}+\cdots+k_{p_{2}}).$
Corollary 2 implies that the values of $k_{2},\ldots,k_{p_{2}}$ depend only on $\ell^{\prime}$ (and $b$), and do not change for different values of $k_{1}=\nu+1$. Since $\ell^{\prime}$ ranges from 0 to $p_{2}-1$, there will be at most $p_{2}$ different sequences of the exponents $k_{2},\ldots,k_{p_{2}}$. If $p_{1}>2p_{2}$, so that $p_{1}-p_{2}>p_{2}$, then these sequences will repeat (as $c$ increases), with different values of $k_{1}$, as illustrated in the following example.
@@
Suppose $p_{2}=7$ and $p_{1}=7a+2$ for some integer $a$. Then for each value of $k_{1}$, there are 7 different apportionment sequences corresponding to $c\in\left[\frac{\ell}{p_{1}-7},\frac{\ell+1}{p_{1}-7}\right)$ where $\ell=\nu 7+\ell^{\prime}=(k_{1}-1)7+\ell^{\prime}$, as shown in Table 2. For each $\ell^{\prime}=0,\ldots,6$, the position of the exponents for which $k_{i}=a+1$ can be determined by evaluating $i_{j}=\left\lceil\frac{7j-\ell^{\prime}}{2}\right\rceil+1$. Thus, when $\ell^{\prime}=0$ (the first row in the table), this occurs when $i=\left\lceil\frac{7(1)-0}{2}\right\rceil+1=5$. When $\ell^{\prime}=1$, this occurs when $i=\left\lceil\frac{7(1)-1}{2}\right\rceil+1=4$, and so on.
@@
Table 1 also includes a row for $c=1$ (D’Hondt’s method). It is the only sequence that does not end in a “1”. In fact, this is true more generally.
@@
Suppose $p_{1}$ and $p_{2}$ are relatively prime and the apportionment sequence for $h=p_{1}+p_{2}$ corresponding to the stationary divisor method with cutpoint $c$ is
$1^{k_{1}},\,2,\,1^{k_{2}},\,2,\,\ldots,1^{k_{j}},\,2,\,1^{k_{p_{2}}},\,2,\,1^{p_{1}-(k_{1}+\cdots+k_{p_{2}})}.$
@@
Substituting $i=p_{2}$ into the inequality (5) in the proof of Proposition 5, we get
@@
If $c=1$, then $\frac{p_{2}}{p_{2}-1+c}=1,$ which implies
@@
Hence, $\sum_{j\leq p_{2}}k_{j}\leq p_{1}<\sum_{j\leq p_{2}}k_{j}+1$, which implies $p_{1}=\sum_{j\leq p_{2}}k_{j}$.
Alternatively, if $p_{1}=\sum_{j\leq p_{2}}k_{j}$, then the left-hand inequality in (7) implies
@@
Note that if $c=1$ and $p_{1}=k_{1}+\cdots+k_{p_{2}}$, then the left-hand inequality in (7) is an equality: $\frac{p_{1}}{p_{1}-1+1}=\frac{p_{2}}{p_{2}-1+1}=1.$ Thus, the last two seats for each party are awarded, first to $p_{1}$ and then to $p_{2}$ using the tie-breaking rule.
@@
Corollary 3.
If $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}=ap_{2}+b$, then the apportionment sequence for $h=p_{1}+p_{2}$ corresponding to D’Hondt’s method is
@@
where $k_{i}=a+1$ if $i\in S=\left\{\lceil\frac{p_{2}}{b}\rceil,\lceil\frac{2p_{2}}{b}\rceil,\ldots,\lceil\frac{(b-1)p_{2}}{b}\rceil,\lceil p_{2}\rceil\right\}$ and $k_{i}=a$ otherwise.
@@
If $p_{1}$ and $p_{2}$ are relatively prime, then there are exactly $p_{1}-p_{2}+1$ different apportionment sequences corresponding to stationary divisor methods.
@@
Suppose $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}>p_{2}+1$ and the apportionment sequence has the repeated pattern of
@@
where the $k_{i}$ satisfy (1) and (2). Then the apportionment sequence is generated by a stationary divisor method with cutpoint $c\in[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}+p_{2}})$ where $\ell=\nu p_{2}+\ell^{\prime}$, $\nu=k_{1}-1$ and $\ell^{\prime}=\max_{i_{j}\in S}\{p_{2}j-(i_{j}-1)b\}$ where $S=\{i_{j}:2\leq i_{j}\leq p_{2}\text{ and }k_{i_{j}}=\lfloor\frac{p_{1}}{p_{2}}\rfloor+1\}$.
We conclude this section by remarking that all the results of this section apply to more general values of $\mathbf{p}=(p_{1},p_{2})$ with $p_{1}$ and $p_{2}$ replaced by $p_{1}/\gcd(p_{1},p_{2})$ and $p_{2}/\gcd(p_{1},p_{2})$ in all expressions for $k_{i}$, etc.